OpenStudy (anonymous):

find any critical numbers of the function. (Enter your answers as a comma-separated list.) g(x) = x5 − 10x3

OpenStudy (anonymous):

You'll have to determine 1st derivative to find critical points: g'(x)=5x^4 - 30x^2 Set g'(x)=0 5x^4 - 30x^2=0 x^4 - 6x^2 = 0 x^2(x^2 - 6) = 0 x^2 = 0 x1=x2=0 x^2 - 6 = 0 x^2 = 6 x3=+sqrt6 x4=-sqrt6 The critical points are the roots of derivative: -sqrt6,0,sqrt6